OpenStudy (anonymous):

What is the y intercept of the line 5x+4y=16

OpenStudy (anonymous):

solve for y first, so 5x + 4y = 16 4y = -5x + 16 y = (-5x + 16)/4 y = (-5x)/4 + 16/4 y = -5x/4 + 4 y = mx + b b = y-intercept b = 4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem